OpenStudy (anonymous):

1. A 12-ft-by-15-ft rectangular swimming pool has a 3-ft-wide no-slip surface around it. What is the outer perimeter of the no-slip surface? (1 point) 78 ft 78 ft2 198 ft 198 ft2

OpenStudy (anonymous):

Do you know where to start with this problem?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do you know how to fin the perimiter?

OpenStudy (anonymous):

yeah you just add all the sides together

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Add 30 and 24, then add 12.

OpenStudy (anonymous):

now you have three more feet on the edges

OpenStudy (anonymous):

@emilyowl your answer isnt even one of the choices

OpenStudy (anonymous):

yeah but what do I do after that?

OpenStudy (anonymous):

add three to all the sides

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so i'm at 66

OpenStudy (anonymous):

yea so am I...

OpenStudy (anonymous):

just give me a sec

OpenStudy (anonymous):

12+3+3 = 18,* 2 sides = 36 15+3+3= 21 * 2 sides = 42 I found this.

OpenStudy (anonymous):

does this make sense to you?

OpenStudy (anonymous):

yeah cuz then u add 36 and 42 to get 78

OpenStudy (anonymous):

yea but do you understand why we got those numbers?

OpenStudy (anonymous):

is it 78 ft or ft^2?

OpenStudy (anonymous):

ft squared is only when you are dealing with area

OpenStudy (anonymous):

yeah cuz 12 and 15 were the swimming pool thanks!

OpenStudy (anonymous):

yep no problem